OpenStudy (zubhanwc3):

3-5 sentences about the chain rule math essayish question.

OpenStudy (zubhanwc3):

I was given the essay: Your friend tells you “I’m never sure when I am supposed to use the chain rule”.Write three to five complete sentences about what you would say to help your friend understand this important concept. i don't really understand the chain rule myself, so im a bit stuck

OpenStudy (anonymous):

ok u use chain rule if f is differentiable at the point u=g(x), and g is differentiable at x, then the composite function (f o g)(x)=f'(g(x)) g'(x) in leibniz notation, if y=f(u) and u=g(x), then dy/dx=dy/du (du/dx), where dy/du is evaluated at u=g(x).

OpenStudy (anonymous):

do need a example

OpenStudy (zubhanwc3):

it doesn't say so.

OpenStudy (anonymous):

so thats a basic rule of chain rule

OpenStudy (zubhanwc3):

alright, thank you very much, i shall give you a medal as your reward.

OpenStudy (anonymous):

ur welcome